# Can Prompt Difficulty be Online Predicted for Accelerating RL Finetuning of Reasoning Models?

### v4

@@ 6 Conclusion
  • Appendix A (Notations Explanation): illustrates the key notations used throughout the paper by walking through one RL finetuning step of reasoning models..
@@ Appendix A Notations Explanation
$p(r_{\tau}^{t,i})=(\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=1]}\cdot(1-\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=0]}\Rightarrow p(\bm{r}^{t}_{\tau}\mid\gamma^{t}_{\tau})=\binom{k}{s^{t}_{\tau}}\cdot(\gamma^{t}_{\tau})^{s^{t}_{\tau}}\cdot(1-\gamma^{t}_{\tau})^{k-s^{t}_{\tau}}\ \text{with}\ s^{t}_{\tau}\triangleq\sum_{j=1}^{k}r_{\tau}^{t,j}.$ (20)
@@
Figure 7: Illustration of the prompt selection, LLM generation, prompt evaluation, and posterior update process during RL finetuning. At each step, a batch of prompts is actively selected based on the predicted success rates $\hat{\gamma}_{\tau}^{t}$ using Thompson Sampling. Then, the model generates multiple responses per prompt and receives binary rewards drawn from a Binomial distribution parameterized by the latent success rate $\gamma_{\tau}^{t}$. These observations are used to update the Beta posterior in a recursive manner.
@@ Appendix B Theoretical Proof
$C_{\tau}^{t}:=\alpha_{\tau}^{t^{\prime}}+\beta_{\tau}^{t^{\prime}}=\lambda C_{\tau}^{t-1}+(1-\lambda)(\alpha_{\tau}^{0}+\beta_{\tau}^{0})+k,\quad t\geq 0.$ (21)
@@
$x_{n}=\frac{1-\lambda^{n}}{1-\lambda}k+x_{-1}.$ (23)
@@
$\bar{\gamma}^{-1}_{\tau}:=\frac{\alpha^{0}_{\tau}}{\alpha^{0}_{\tau}+\beta^{0}_{\tau}}$ (25)
@@
$\bar{\gamma}^{t}_{\tau}$ $=\frac{\lambda\alpha^{t}_{\tau}+(1-\lambda)\alpha_{\tau}^{0}+s^{t}_{\tau}}{C_{\tau}^{t}}$ (26)
@@
$\bar{\gamma}^{t}_{\tau}=w_{1}\cdot\bar{\gamma}^{t-1}_{\tau}+w_{2}\cdot\bar{\gamma}_{\tau}^{-1}+w_{3}\cdot\bar{s}^{t}_{\tau}$ (27)
@@
$|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|\leq w_{1}\cdot|\bar{\gamma}^{t-1}_{\tau}-\gamma^{t}_{\tau}|+w_{2}\cdot|\bar{\gamma}_{\tau}^{-1}-\gamma^{t}_{\tau}|+w_{3}\cdot|\bar{s}^{t}_{\tau}-\gamma^{t}_{\tau}|.$ (28)
@@
$|\bar{\gamma}^{t-1}_{\tau}-\gamma^{t}_{\tau}|\leq|\bar{\gamma}^{t-1}_{\tau}-\gamma^{t-1}_{\tau}|+|\gamma^{t-1}_{\tau}-\gamma^{t}_{\tau}|\leq\epsilon_{t-1}+\delta.$ (29)
@@
$\mathbb{P}\left(|\bar{s}^{t}_{\tau}-\gamma^{t}_{\tau}|\geq\eta\right)\leq 2\exp(-2k\eta^{2}).$ (31)
@@
$\epsilon_{t}=|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|\leq w_{1}\cdot(\epsilon_{t-1}+\delta)+w_{2}\cdot 1+w_{3}\cdot\eta$ (32)
@@
$w_{1}=\lambda\cdot\frac{C_{\tau}^{t-1}}{C_{\tau}^{t}}\leq\lambda.$ (33)
@@
$w_{2}=(1-\lambda)\cdot\frac{(\alpha_{\tau}^{0}+\beta_{\tau}^{0})}{C_{\tau}^{t}}\leq(1-\lambda)\cdot\frac{C_{\tau}^{0}}{C_{\tau}^{1}}\leq\frac{1-\lambda}{2}.$ (34)
@@
$|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|<\lambda\cdot(\epsilon_{t-1}+\delta)+\frac{(1-\lambda)}{2}+\eta.$ (35)
@@
$|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|\lessapprox\epsilon_{t-1}+\eta.$ (36)
@@ C.1.2 Planning
Evaluation is conducted on two benchmarks: a 512-problem held-out split from Countdown-34 (CD-34), and a 512-problem subset from Countdown-4 (CD-4), a more challenging variant from https://huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-4. Unlike CD-34, CD-4 provides 4 source numbers per problem, which significantly increases the search space and problem difficulty. Training curves are plotted using CD-34.
@@
Following Pan et al. (2025), we include a format term in the reward function:
@@ C.3 Training Details
At each training step, $k=8$ responses per prompt are sampled to estimate advantages, using temperature 1.0 and $\texttt{top\_p}=1.0$. Evaluation is based on pass@1, computed from 16 independent generations per prompt with temperature 0.6 and $\texttt{top\_p}=0.95$, following Luo et al. (2025b). We disable the KL penalty ($\beta=0$) for MATH and Countdown, following Yu et al. (2025), but retain it for Geometry3k with $\beta=0.1$ (3B) and $\beta=0.3$ (7B) to ensure training stability. Training batch sizes are set to 256 for MATH and Countdown with mini-batch sizes of 128 and 64, repectively, and 512 for Geometry3k with a mini-batch size of 256. The maximum response length is 8192 for MATH, and 1024 for Countdown and Geometry3k. Entropy regularization is applied with coefficient 0.001, following Luo et al. (2025b). Optimization is performed using Adam (Kingma & Ba, 2014) with a learning rate of $1\mathrm{e}{-6}$, beta $(0.9,0.999)$, no warm-up, and weight decay $0.01$. The Clip-Higher strategy from Yu et al. (2025) is applied, which decouples clipping ranges, with $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$. All experiments are conducted on 8 NVIDIA A100 or H100 80GB GPUs.
@@
We set the Beta prior as $(\alpha^{0},\beta^{0})=(1,1)$, target success probability $\gamma^{*}=0.5$ for all training tasks. The decay factor $\lambda$ is set to $0.5$ for Countdown and Geometry3k, $1$ for MATH. Since the primary objective of this work is to address the two research questions, and the current performance is satisfactory, we do not emphasize performance optimization. We leave extensive hyperparameter tuning for future work.
@@
Appendix D Additional Results
D.1 Evaluation Results
Table 3: Evaluation results on Countdown. Models trained with different prompt selection methods are evaluated on two benchmarks: Countdown-34 (CD-34) and Countdown-4 (CD-4). MoPPS consistently outperforms Uniform without requiring any additional LLM inference, and matches or surpasses DS while using substantially fewer LLM-generated rollouts.
@@ D.1 Evaluation Results
Table 4: Evaluation results on Geometry. Models trained with different prompt selection methods are evaluated on the test split of Geometry3k (Geo3k test). MoPPS consistently outperforms Uniform without requiring any additional LLM inference, and surpasses DS while using substantially fewer LLM-generated rollouts.
@@
Table 5: Evaluation across mathematics benchmarks with maximum response length 32,768. ‘+’ indicates finetuning with the corresponding method. Accuracy is computed as the average pass@1 over 16 independent generations per problem. ‘Avg.’ denotes average accuracy across benchmarks, and ‘Rollouts’ indicates the number of rollout samples during finetuning. Bold indicates the best result; underlined indicates the second best. ‘MoPPS w/ prior’ means incorporating prior knowledge into MoPPS.
@@
D.2 Other Analysis
D.2.1 Algorithm Compatibility.
@@ D.2 Other Analysis
Figure 8: Training curves on Countdown with PPO (both $k=1$ and $k=8$) and Reinforce++. MoPPS consistently accelerates convergence and improves final accuracy compared to uniform prompt selection across different RL algorithms and sampling number settings. The Spearman rank correlation further shows that MoPPS reliably predicts prompt difficulty across varied setups.
D.2.2 Rollout Efficiency.
@@
Figure 9: Training curves plotted against the number of rollouts generated by LLM during training. MoPPS achieves comparable accuracy with significantly fewer rollouts than DS.
Figure 10: Ablation study on target success rate $\gamma^{*}$. (a) Performance comparison under different $\gamma^{*}$ values on the Countdown task with Qwen2.5-3B. (b) Spearman rank correlation under different $\gamma^{*}$ values. (c,d) Number of ineffective prompts (i.e., Solve-All or Solve-None) per batch. These results support choosing intermediate success rates for stronger learning signals.
D.2.3 Ablation Study on Target Success Rate.
@@
D.2.4 Ablation Study on Temporal Discounting.
@@
Figure 11: Ablation study on the temporal discounting strategy. We evaluate MoPPS under different $\lambda$ values, including disabling temporal discounting (MoPPS ($\lambda=1$)) and using only current feedback (MoPPS ($\lambda=0$)).
@@ Discussion.
Notably, the primary goal of this work is to demonstrate that prompt difficulty can be online predicted and leveraged for accelerated RL finetuning via appropriate selection strategies. As such, the focus lies not on absolute performance tuning, but on validating the effectiveness of our predictive selection framework. Even so, MoPPS already achieves strong and competitive results, matching Dynamic Sampling (DS) while requiring significantly fewer LLM rollouts. The ablation on temporal discounting further supports the soundness of our design. We therefore leave finer hyperparameter tuning, e.g., the decay factor $\lambda$ and target success rate $\gamma^{*}$ to future work. Moreover, in tasks like MATH where training dynamics are relatively stable, we disable discounting to simplify the implementation and tuning requirements.
D.2.5 Reduction of Ineffective Prompts.
Figure 12: Number of ineffective prompts (i.e., Solve-All or Solve-None) per batch. MoPPS substantially reduces such prompts compared to uniform sampling, leading to more efficient training.
D.2.6 Response Length.
@@
Figure 13: Mean response length during MATH training. MoPPS and DS both elicit longer responses than Uniform, explaining improved performance.
D.2.7 Selected Prompt Length.
@@
Figure 14: Mean prompt length during MATH training. MoPPS and DS tend to select longer prompts than Uniform.
@@ Appendix E Data Examples
The prompt templates for MATH and Geometry3k are adopted from the verl framework, while the template for Countdown follows the format introduced in Pan et al. (2025).

### v5

@@ 6. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (NSFC) with the Number # 62306326, the National Key R&D Program of China under Grant 2018AAA0102801, and the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (JYB2025XDXM503). We thank all reviewers for their insightful comments and constructive suggestions.
@@ Acknowledgements.
Figure 7. Illustration of the prompt selection, LLM generation, prompt evaluation, and posterior update process during RL finetuning. At each step, a batch of prompts is actively selected based on the predicted success rates $\hat{\gamma}_{\tau}^{t}$ using Thompson Sampling. Then, the model generates multiple responses per prompt and receives binary rewards drawn from a Binomial distribution parameterized by the latent success rate $\gamma_{\tau}^{t}$. These observations are used to update the Beta posterior in a recursive manner.
@@
  • Appendix A (Notations Explanation): illustrates the key notations used throughout the paper by walking through one RL finetuning step of reasoning models.
@@ Appendix A Notations Explanation
(20) $p(r_{\tau}^{t,i})=(\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=1]}\cdot(1-\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=0]}\Rightarrow$
$p(\bm{r}^{t}_{\tau}\mid\gamma^{t}_{\tau})=\binom{k}{s^{t}_{\tau}}\cdot(\gamma^{t}_{\tau})^{s^{t}_{\tau}}\cdot(1-\gamma^{t}_{\tau})^{k-s^{t}_{\tau}}\ \text{with}\ s^{t}_{\tau}\triangleq\sum_{j=1}^{k}r_{\tau}^{t,j}.$
@@ Appendix B Theoretical Proof
(21) $C_{\tau}^{t}:=\alpha_{\tau}^{t^{\prime}}+\beta_{\tau}^{t^{\prime}}=\lambda C_{\tau}^{t-1}+(1-\lambda)(\alpha_{\tau}^{0}+\beta_{\tau}^{0})+k,\quad t\geq 0.$
@@
(23) $x_{n}=\frac{1-\lambda^{n}}{1-\lambda}k+x_{-1}.$
@@
(25) $\bar{\gamma}^{-1}_{\tau}:=\frac{\alpha^{0}_{\tau}}{\alpha^{0}_{\tau}+\beta^{0}_{\tau}}$
@@
(26) $\bar{\gamma}^{t}_{\tau}$ $=\frac{\lambda\alpha^{t}_{\tau}+(1-\lambda)\alpha_{\tau}^{0}+s^{t}_{\tau}}{C_{\tau}^{t}}$
@@
(27) $\bar{\gamma}^{t}_{\tau}=w_{1}\cdot\bar{\gamma}^{t-1}_{\tau}+w_{2}\cdot\bar{\gamma}_{\tau}^{-1}+w_{3}\cdot\bar{s}^{t}_{\tau}$
@@
(28) $|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|\leq w_{1}\cdot|\bar{\gamma}^{t-1}_{\tau}-\gamma^{t}_{\tau}|+w_{2}\cdot|\bar{\gamma}_{\tau}^{-1}-\gamma^{t}_{\tau}|+w_{3}\cdot|\bar{s}^{t}_{\tau}-\gamma^{t}_{\tau}|.$
@@
(29) $|\bar{\gamma}^{t-1}_{\tau}-\gamma^{t}_{\tau}|\leq|\bar{\gamma}^{t-1}_{\tau}-\gamma^{t-1}_{\tau}|+|\gamma^{t-1}_{\tau}-\gamma^{t}_{\tau}|\leq\epsilon_{t-1}+\delta.$
@@
(31) $\mathbb{P}\left(|\bar{s}^{t}_{\tau}-\gamma^{t}_{\tau}|\geq\eta\right)\leq 2\exp(-2k\eta^{2}).$
@@
(32) $\epsilon_{t}=|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|\leq w_{1}\cdot(\epsilon_{t-1}+\delta)+w_{2}\cdot 1+w_{3}\cdot\eta$
@@
(33) $w_{1}=\lambda\cdot\frac{C_{\tau}^{t-1}}{C_{\tau}^{t}}\leq\lambda.$
@@
(34) $w_{2}=(1-\lambda)\cdot\frac{(\alpha_{\tau}^{0}+\beta_{\tau}^{0})}{C_{\tau}^{t}}\leq(1-\lambda)\cdot\frac{C_{\tau}^{0}}{C_{\tau}^{1}}\leq\frac{1-\lambda}{2}.$
@@
(35) $|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|<\lambda\cdot(\epsilon_{t-1}+\delta)+\frac{(1-\lambda)}{2}+\eta.$
@@
(36) $|\bar{\gamma}^{t}_{\tau}-\gamma^{t}_{\tau}|\lessapprox\epsilon_{t-1}+\eta.$
@@ C.1.2. Planning
Evaluation is conducted on two benchmarks: a 512-problem held-out split from Countdown-34 (CD-34), and a 512-problem subset from Countdown-4 (CD-4), a more challenging variant from huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-4. Unlike CD-34, CD-4 provides 4 source numbers per problem, which significantly increases the search space and problem difficulty. Training curves are plotted using CD-34.
@@
Following (Pan et al., 2025), we include a format term in the reward function:
@@ C.3. Training Details
At each training step, $k=8$ responses per prompt are sampled to estimate advantages, using temperature 1.0 and $\texttt{top\_p}=1.0$. Evaluation is based on pass@1, computed from 16 independent generations per prompt with temperature 0.6 and $\texttt{top\_p}=0.95$, following Luo et al. (2025b). We disable the KL penalty ($\beta=0$) for MATH and Countdown, following Yu et al. (2025), but retain it for Geometry3k with $\beta=0.1$ (3B) and $\beta=0.3$ (7B) to ensure training stability. Training batch sizes $\mathcal{B}$ are set to 256 for MATH and Countdown with mini-batch sizes of 128 and 64, repectively, and 512 for Geometry3k with a mini-batch size of 256. The maximum response length is 8192 for MATH, and 1024 for Countdown and Geometry3k. Entropy regularization is applied with coefficient 0.001, following Luo et al. (2025b). Optimization is performed using Adam (Kingma and Ba, 2014) with a learning rate of $1\mathrm{e}{-6}$, beta $(0.9,0.999)$, no warm-up, and weight decay $0.01$. The Clip-Higher strategy from Yu et al. (2025) is applied, which decouples clipping ranges, with $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$. All experiments are conducted on 8 NVIDIA A100 or H100 80GB GPUs.
@@
We set the Beta prior as $(\alpha^{0},\beta^{0})=(1,1)$, target success probability $\gamma^{*}=0.5$ for all training tasks. The decay factor $\lambda$ is set to $0.5$ for Countdown and Geometry3k, $1$ for MATH. Since the primary objective of this work is to address the two research questions, and the current performance is satisfactory, we do not emphasize performance optimization. We leave extensive hyperparameter tuning for future work. The candidate batch size $\hat{\mathcal{B}}$ is set to $16\times\mathcal{B}$, which is relatively large. For Countdown and Geometry3k, this means the candidate prompt batch $\mathcal{T}^{\hat{\mathcal{B}}}$ effectively covers the entire pool $\mathcal{T}$.
@@
Table 3. Evaluation results on Countdown. Models trained with different prompt selection methods are evaluated on two benchmarks: Countdown-34 (CD-34) and Countdown-4 (CD-4). MoPPS consistently outperforms Uniform without requiring any additional LLM inference, and matches or surpasses DS while using substantially fewer rollouts.
@@
Table 4. Evaluation results on Geometry. Models trained with different prompt selection methods are evaluated on the test split of Geometry3k (Geo3k test). MoPPS consistently outperforms Uniform without requiring any additional LLM inference, and surpasses DS while using substantially fewer LLM-generated rollouts.
@@
Table 5. Evaluation across mathematics benchmarks with maximum response length 32,768. ‘+’ indicates finetuning with the corresponding method. Accuracy is computed as the average pass@1 over 16 independent generations per problem. ‘Avg.’ denotes average accuracy across benchmarks, and ‘Rollouts’ indicates the number of rollout samples during finetuning. Bold indicates the best result; underlined indicates the second best. ‘MoPPS w/ prior’ means incorporating prior knowledge.
@@
Appendix D Additional Results
D.1. Evaluation Results
D.2. Other Analysis
Figure 8. Training curves on Countdown with PPO (both $k=1$ and $k=8$) and Reinforce++. MoPPS consistently accelerates convergence and improves final accuracy compared to uniform prompt selection across different RL algorithms and sampling number settings. The Spearman rank correlation further shows that MoPPS reliably predicts prompt difficulty across varied setups.
D.2.1. Algorithm Compatibility.
@@ D.2. Other Analysis
Figure 9. Training curves plotted against the number of rollouts generated by LLM during training. MoPPS achieves comparable accuracy with significantly fewer rollouts than DS.
D.2.2. Rollout Efficiency.
@@
Figure 10. Number of ineffective prompts (i.e., Solve-All or Solve-None) per batch. MoPPS substantially reduces such prompts compared to uniform sampling, leading to more efficient training.
D.2.3. Reduction of Ineffective Prompts.
Figure 11. Ablation study on target success rate $\gamma^{*}$. (a) Performance comparison under different $\gamma^{*}$ values on the Countdown task with Qwen2.5-3B. (b) Spearman rank correlation under different $\gamma^{*}$ values. (c,d) Number of ineffective prompts (i.e., Solve-All or Solve-None) per batch. These results support choosing intermediate success rates for stronger learning signals.
D.2.4. Ablation Study on Target Success Rate.
@@
Figure 12. Ablation study on the temporal discounting strategy. We evaluate MoPPS under different $\lambda$ values, including disabling temporal discounting (MoPPS ($\lambda=1$)) and using only current feedback (MoPPS ($\lambda=0$)).
D.2.5. Ablation Study on Temporal Discounting.
@@ Discussion.
Notably, the primary goal of this work is to demonstrate that prompt difficulty can be online predicted and leveraged for accelerated RL finetuning via appropriate selection strategies. As such, the focus lies not on absolute performance tuning, but on validating the effectiveness of our predictive selection framework. Even so, MoPPS already achieves strong and competitive results, matching Dynamic Sampling (DS) while requiring significantly fewer LLM rollouts. The ablation studies further supports the soundness of the design and the proposed component. We therefore leave finer hyperparameter tuning, e.g., the decay factor $\lambda$ and target success rate $\gamma^{*}$ to future work. Moreover, in tasks like MATH where training dynamics are relatively stable, we disable discounting to simplify the implementation and tuning requirements.
D.2.6. Ablation Study on Candidate Batch Size.
We evaluate the effect of candidate batch size $\hat{\mathcal{B}}$ on Countdown, as shown in Table 6. The results show that performance improves as the candidate size increases, due to the expanded exploration space. This is a key advantage of MoPPS: prior methods relying on real evaluations are limited by computational cost and cannot scale, whereas MoPPS uses lightweight prediction models to efficiently explore a much larger set of prompts. In practice, setting $\hat{\mathcal{B}}$ to the entire prompt pool is viable.
Table 6. Performance with Different Candidate Batch Sizes on Countdown.
Candidate Size $\hat{\mathcal{B}}$ 3B CD-34 3B CD-4 7B CD-34 7B CD-4
$1\times\mathcal{B}$ (Uniform) 69.87 39.42 77.84 53.27
$2\times\mathcal{B}$ 72.29 43.15 78.53 54.40
$4\times\mathcal{B}$ 73.17 45.70 79.86 58.04
$\approx 8\times\mathcal{B}$ (Entire Pool) 74.46 47.02 82.18 59.16
D.2.7. Response Length.
@@
Figure 13. Mean response length during MATH training. MoPPS and DS both elicit longer responses than Uniform, explaining improved performance.
D.2.8. Selected Prompt Length.
@@
Figure 14. Mean prompt length during MATH training. MoPPS and DS tend to select longer prompts than Uniform.
@@ Appendix E Data Examples
The prompt templates for MATH and Geometry3k are adopted from the official verl framework, while the template for Countdown follows the format introduced in (Pan et al., 2025).